OpenStudy (anonymous):

For the function y=f(x), what is the ordered pair for the point on the graph when x=2m+3? (x, f(m)) (x, 2m+3) (2m+3, f(m)) (2m+3, f(2m+3))

OpenStudy (anonymous):

If x = 2m+3, what is the x-coordinate?

OpenStudy (anonymous):

The answer to the equation?

OpenStudy (anonymous):

If you read the question, it's talking about a function that has an ordered pair. And x = 2m+3. You want to write the ordered pair. So an ordered pair has an x-coordinate and a y-coordinate. What will be the x-coordinate?

OpenStudy (anonymous):

m? (I'm not very good at math)

OpenStudy (anonymous):

It says that x = 2m+ 3, so the x-coordinate is 2m + 3. What is the y-ccordinate?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

no. y = f(x)...do you know what that means?

OpenStudy (anonymous):

That makes the equation a function?

OpenStudy (anonymous):

Do you understand what "y = f(x)" means?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

well, that's part of the question. You have to go back to understand what "y = f(x)" means; then you will be able to do this question.

OpenStudy (anonymous):

The best thing is to restudy what a function is, and what y = f(x) means. You can post a new question here from an earlier lesson on functions.

OpenStudy (anonymous):

@amistre64 Can you please walk me through this question?

OpenStudy (anonymous):

@jim_thompson5910 Can you help me?

jimthompson5910 (jim_thompson5910):

y = f(x) ... given y = f(2m+3) ... replace x with 2m+3 (since x = 2m+3) So we know that x = 2m+3 and y = f(2m+3) which means what?

OpenStudy (anonymous):

Which means that we'll get the x coordinate(s) by doing the x = given equation, and that we'll get the y coordinate(s) by doing the y = given equation?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

well replace "equation" with "expression" and you're more correct so what is the answer?

OpenStudy (anonymous):

d?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

:D Thank you so much. You think you can help me with some more questions? It'd be great if you can!

jimthompson5910 (jim_thompson5910):

sure maybe 2 more

OpenStudy (anonymous):

Okay, thanks a lot!